OpenStudy (anonymous):

evaluate the integral (dx/(2root(x-4)+x))

OpenStudy (anonymous):

\[\int\limits_{ }^{ }\left(\begin{matrix}dx \\ 2\sqrt{x-4}+x\end{matrix}\right)\]

OpenStudy (anonymous):

\[I=\int\limits \frac{ dx }{ 2\sqrt{x-4}+x}\] rationalise the denominator

OpenStudy (anonymous):

you got it.Then show me your work.

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }\int\limits_{ }^{}\frac{ \sqrt{x-4}-x }{ x^2-x-4 } dx\]

OpenStudy (anonymous):

there is a - sign i missed

OpenStudy (anonymous):

-x^2+x is what it should be on the bottom

OpenStudy (anonymous):

no .it is going wrong way. put \[\sqrt{x-4}=t,x-4=t^2,x=t^2+4,dx=2t~dt\]

OpenStudy (anonymous):

\[I=\int\limits \frac{ 2t~dt }{ 2t+t^2+4 }=\int\limits \frac{ 2t+2-2 }{t^2+2t+4 }dt\] can you go further?

OpenStudy (anonymous):

how do you get 2t on the denominator?

OpenStudy (anonymous):

\[2\sqrt{x-4}=2t\]

OpenStudy (anonymous):

ok, I had pulled out the 2 incorectly

OpenStudy (anonymous):

int (2t+2)/(t+2)^2 -2/(t+2)^2 dt

OpenStudy (anonymous):

\[=\int\limits \frac{ 2t+2 }{t^2+2t+4 }dt-2 \int\limits \frac{ dt }{(t^2+2t+1)+3 }\]

OpenStudy (anonymous):

\[=\ln \left| t^2+2t+4 \right|+?\]

OpenStudy (anonymous):

is there a substitution im not seeing? i dont know where you get that

OpenStudy (anonymous):

\[\int\limits \frac{ dx }{ x^2+a^2 }=\frac{ 1 }{ a }\tan^{-1} \frac{ x }{ a }\]

OpenStudy (anonymous):

the ln(*) part

OpenStudy (anonymous):

\[\int\limits \frac{ f' \left( x \right) }{ f \left( x \right) }dx=\ln \left| f \left( x \right) \right|\]

OpenStudy (anonymous):

\[\frac{ d }{ dt }\left( t^2+2t+4 \right)=\frac{ d }{ dt }(t^2) +\frac{ d }{ dt }(2t)+\frac{ d }{dt }( 4)=2t+2+0\]

OpenStudy (anonymous):

\[\int\limits \frac{ dt }{ \left( t+1 \right)^2+\left( \sqrt{3} \right)^2 }\]

OpenStudy (anonymous):

calculate it . formula is given above.

OpenStudy (anonymous):

now you got it or have some quiry.